### v1

@@ Appendix D Further Numerical Experiments
First, we discuss the specifics of the construction of our test matrices. Recall that each of our test matrices is an $n\times n$ PSD matrix ${\mathbf{A}}$ with $n=4096$, and our task is solving a linear system $({\mathbf{A}}+\phi{\mathbf{I}}){\mathbf{x}}={\mathbf{b}}$ where we choose $\phi=0.001$ and generate ${\mathbf{b}}$ from the standard normal distribution. Note that we choose $n$ to be a power of $2$ simply for the convenience of implementing randomized Hadamard transform (RHT). In general this is not necessary, since we can still implement it by finding the closest power of $2$ larger than $n$, enlarging the matrix to that dimension by padding with 00 entries, and truncating back to the original dimension at the end.
We consider the following two classes of test matrices.
@@
To evaluate our algorithm in a practical setting that naturally exhibits large outlying eigenvalues, we consider applying kernel transformation to four real-world datasets (Abalone and Phoneme from OpenML [VvRBT13], California_housing and Covtype from Scikit-learn [PVG+11]). We truncate each dataset to its first $n$ rows to get matrix ${\mathbf{\Phi}}\in{\mathds{R}}^{n\times m}$. For $(i,j)\in[n]\times[n]$, we define the kernel matrix ${\mathbf{A}}\in\mathcal{S}_{d}^{+}$ so that ${\mathbf{A}}_{ij}=\mathcal{K}({\mathbf{\Phi}}_{i},{\mathbf{\Phi}}_{j})$ where ${\mathbf{\Phi}}_{i},{\mathbf{\Phi}}_{j}$ are the $i$-th and $j$-th rows of ${\mathbf{\Phi}}$, respectively, and $\mathcal{K}$ is a kernel function. We consider two types of kernel functions: Gaussian, $\mathcal{K}({\mathbf{\Phi}}_{i},{\mathbf{\Phi}}_{j})=\exp(-\gamma\|{\mathbf{\Phi}}_{i}-{\mathbf{\Phi}}_{j}\|^{2})$, and Laplacian, $\mathcal{K}({\mathbf{\Phi}}_{i},{\mathbf{\Phi}}_{j})=\exp(-\gamma\|{\mathbf{\Phi}}_{i}-{\mathbf{\Phi}}_{j}\|)$. For both choices, we set the width parameter $\gamma$ among two values, $\{0.1,0.01\}$. This leads to four different test matrices for each of the four datasets, giving a total of 16 matrices.
D.1 Testing Acceleration and Block Memoization
In this section we test the effect of both the adaptive acceleration scheme as well as the block memoization technique used in our methods (see Figure 1 in Section 6 and Figure 3 below). For all tasks, we used variants of CD++ (Algorithm 3), since it is designed specifically for solving PSD linear systems. Recall that CD++ can also be viewed as the dual algorithm of Kaczmarz++ (Algorithm 1) in the sense that instead of sampling principal submatrices of ${\mathbf{A}}$, it is sampling rows of matrix ${\mathbf{\Phi}}$. So, in that sense, these experiments simultaneously provide an evaluation of the convergence properties of Kaczmarz++ (Algorithm 1). We consider four variants of our method, as shown in Table 1, to identify the effect of each component individually. For clarity, we explain their differences below.
@@ D.1 Testing Acceleration and Block Memoization
  • CD: The classical block coordinate descent method, without acceleration or block memoization, but still preprocessed with RHT;
  • CD+Accel: Coordinate descent with adaptive acceleration, but without block memoization, i.e., Algorithm 3 modified so that in line 6 we always choose to sample a new block $S\sim{[n]\choose s}$, and hence, do not save the Cholesky factors;
@@
  • CD+Memo: Coordinate descent with block memoization (but without adaptive acceleration), i.e., Algorithm 3 without lines 14-20, and with $\eta=0$ (as opposed to $\eta=\frac{s}{2n}$), meaning that we no longer maintain the adaptive momentum term ${\mathbf{m}}_{t}$ from line 12;
@@
  • Full CD++: Algorithm 3 as given in the pseudocode, equipped with both adaptive acceleration and block memoization.
Throughout this section, we set the sketch size $s$ to be $\{25,50,100,200\}$, and measure the convergence through the residual error defined as $\epsilon_{t}\coloneqq\|{\mathbf{A}}_{\phi}{\mathbf{x}}_{t}-{\mathbf{b}}\|/\|{\mathbf{b}}\|$ where ${\mathbf{A}}_{\phi}\coloneqq{\mathbf{A}}+\phi{\mathbf{I}}$. For each plot we run all methods $10$ times and take an average.
@@
By comparing the curves, we can see that in all cases, adding block memoization slightly worsens the convergence rate (by comparing CD and CD+Memo, or CD+Accel and Full CD++). This is actually suggested by theory, since with block memoization we are essentially not sampling all possible blocks of coordinates, thus reducing the quality of the block sampling distribution. However it is noteworthy that this phenomenon is very insignificant especially when we compare CD+Accel with Full CD++, which suggests that our online block selection scheme (described in Section 5.2) works well, and that $\tilde{O}(n/s)$ blocks are sufficient for fast convergence, matching our theory.
@@
From the plots, we can also see that the adaptive acceleration plays an important role for both CD and CD with block memoization, showing a comparable improvement for both cases. We also observe that the effect of adaptive acceleration is more significant when the sketch size $s$ is smaller - this also aligns with the theory, since for smaller $s$ the tail Demmel condition number $\bar{\kappa}_{s}$ is larger, thus the effect of acceleration reducing the dependence on $\bar{\kappa}_{s}$ from second to first power is also more significant.
To conclude, the experiments in this section suggest that both CD with and without block memoization benefit from our adaptive acceleration scheme. Based on this, in next section we will carry out more experiments testing the effect of block memoization; namely, we will test the performance of CD+Accel and Full CD++ alongside Krylov-type methods including CG and GMRES.
Figure 3: Convergence plots for different variants of CD++ (see Table 1) using four block sizes (continuation of Figure 1). Check https://github.com/EdwinYang7/kaczmarz-plusplus for plots on remaining real-world datasets.
D.2 Comparison with Krylov Subspace Methods
@@
Next, we test the convergence of our Full CD++ alongside CD+Accel, against Krylov-type methods, conjugate gradient (CG) and GMRES. In this experiment we count the FLOPs it takes to converge for each method, to showcase the advantages of our methods compared to Krylov-type methods.
@@
Figure 4: Computational cost comparison, measuring floating point operations (FLOPs) against the normalized residual error (6.1) for Full CD++, alongside baselines CD+Accel, CG, and GMRES (continuation of Figure 2).
@@
For the FLOPs of CG, we look into the source code from scipy.sparse.linalg and approximate it by $2n^{2}+11n$ per iteration. For the FLOPs counting of GMRES, we look into the source code from pyamg.krylov and approximate it by $2n^{2}T+4nT(T+1)$, where $T$ is the number of iterations. For the FLOPs counting of our methods (CD+Accel, Full CD++), we maintain a counter for FLOPs for each iteration, which includes both the cases of using (Full CD++) and not using (CD+Accel) block memoization. Specifically, for Full CD++ which leverages block memoization, if a new block is sampled, then we do the Cholesky factorization and increase the FLOPs by $s^{3}/3$, where $s$ is the block size; if we sample from the already sampled set of blocks ${\mathcal{B}}$, then Cholesky factorization is not needed and this cost is omitted. Notice that for CD+Accel and Full CD++, there is a pre-processing step of applying the RHT which takes extra FLOPs. We count them by following Appendix C, and this is reflected in the plots, since the convergence curves of our methods are shifted by the cost of the RHT. Here, we take advantage of our fast SymFHT implementation (Algorithm 2), which reduces that cost by half. Throughout the section, we choose the block size $s=200$, which tends to work well for the size of our test matrices. Further optimizing the block size, or choosing it dynamically, is an interesting direction for future work. For each plot, we run CD+Accel and Full CD++ $5$ times and take average to reduce the noise.
@@
From the experiments (Figure 4), we can see that CG converges very slowly and is not comparable with other methods. By comparing CD+Accel and Full CD++, we can see that the block memoization technique gives a significant improvement in FLOPs, since Full CD++ successfully reduces the expensive Cholesky factorization step. This improvement is more significant if we are running more iterations (i.e., aiming for higher accuracy). By comparing Full CD++ and GMRES, we can see that in most cases Full CD++ performs better in the “low to medium accuracy level”, while GMRES sometimes beats Full CD++ in the “high accuracy level”.
@@
For example, for the abalone dataset with Gaussian kernel with width=$0.01$, GMRES starts to perform better after the residual reaches $10^{-7}$. However, we note that for abalone, phoneme and california_housing dataset with Gaussian kernel with width=$0.1$, or with Laplacian kernel with width=$0.01$, our Full CD++ outperforms GMRES even when the residual reaches $10^{-12}$. These phenomena are dependent on the spectrum (especially the top eigenvalues) of matrix ${\mathbf{A}}+\phi{\mathbf{I}}$, which is reflected in the different choices of kernel and width.
Problem CG GMRES CD++
Dataset Kernel Width 1e-4 1e-8 1e-4 1e-8 1e-4 1e-8
Abalone Gaussian 0.1 7.63e9 3.06e10 1.47e9 5.17e9 4.64e8 3.26e9
0.01 1.88e9 6.89e9 8.50e8 1.75e9 2.97e8 2.11e9
Laplacian 0.1 8.64e9 3.22e10 2.72e9 8.14e9 2.22e9 8.13e9
0.01 1.11e9 1.37e10 5.41e8 4.11e9 2.40e8 3.09e9
Phoneme Gaussian 0.1 7.80e9 3.29e10 1.79e9 4.98e9 4.86e8 3.23e9
0.01 3.70e8 4.23e9 3.37e8 1.33e9 2.23e8 1.80e9
Laplacian 0.1 7.22e9 3.73e10 2.29e9 8.46e9 1.65e9 8.96e9
0.01 2.49e9 1.41e10 7.81e8 3.97e9 2.80e8 3.10e9
California Housing Gaussian 0.1 2.45e9 3.27e10 1.02e9 6.14e9 2.40e8 3.88e9
0.01 2.15e9 9.51e9 5.75e8 2.18e9 2.27e8 1.99e9
Laplacian 0.1 5.44e9 2.34e10 2.04e9 6.85e9 1.52e9 6.39e9
0.01 1.31e9 1.36e10 5.75e8 4.15e9 2.40e8 3.06e9
Covtype Gaussian 0.1 1.65e10 5.55e10 4.87e9 1.70e10 5.71e9 3.64e10
0.01 9.31e9 3.87e10 2.00e9 7.69e9 9.78e8 5.66e9
Laplacian 0.1 6.79e9 2.24e10 3.23e9 8.34e9 2.85e9 1.03e10
0.01 3.60e9 2.37e10 1.23e9 5.95e9 4.60e8 4.56e9
Synthetic Low-Rank Effective rank $=25$ 1.68e9 3.26e9 1.44e9 1.83e9 1.31e9 2.79e9
Effective rank $=50$ 2.59e9 5.17e9 2.43e9 3.26e9 1.53e9 3.21e9
Effective rank $=100$ 3.16e9 6.89e9 2.65e9 5.67e9 1.91e9 3.89e9
Effective rank $=200$ 3.26e9 8.00e9 2.75e9 8.34e9 2.92e9 6.10e9
@@
Table 2: Comparison of the FLOPs needed to achieve given error threshold, $\epsilon\in\{10^{-4},10^{-8}\}$, for different algorithms. Bold values indicate the best performance for a given error threshold.
@@
In Table 2 we show the FLOPs it takes for different methods (CG, GMRES, CD+Accel and Full CD++) to achieve the given accuracy in detail. Here we set $\epsilon=10^{-4}$ as the mid-level accuracy and $\epsilon=10^{-8}$ as the high-level accuracy. For mid-level accuracy, we can see that Full CD++ outperforms GMRES in 18 out of 20 tasks, showing that our method converges super fast at early stages. For high-level accuracy, we can see that Full CD++ still outperforms GMRES at 14 out of 20 tasks.
@@
To conclude, the experiments in this section show that in the measurement of FLOPs, our CD++ outperforms Krylov methods including CG and GMRES in almost all mid-level accuracy tasks, as well as most high-level accuracy tasks.
D.3 Testing Regularization in Projection

### v2

@@ Appendix F Further Numerical Experiments
First, we discuss the specifics of the construction of our test matrices. We consider two classes of test matrices.
@@
For the task of testing Kaczmarz++ (Algorithm 5), each of the test matrices is an $m\times n$ rectangular matrix ${\mathbf{A}}$ with $m=4096$ and $n=1024$ generated as above, and our task is solving a linear system ${\mathbf{A}}{\mathbf{x}}={\mathbf{b}}$ where ${\mathbf{b}}$ is generated from the standard normal distribution. On the other hand, for the task of evaluating CD++, we use make_low_rank_matrix to construct a $\times 4096$ 4096 matrix ${\mathbf{\Phi}}$, and then compute ${\mathbf{A}}={\mathbf{\Phi}}{\mathbf{\Phi}}^{\scriptscriptstyle{\top}}$. We then solve a PSD linear system $({\mathbf{A}}+\phi{\mathbf{I}}){\mathbf{x}}={\mathbf{b}}$, with standard normal ${\mathbf{b}}$ and $\phi=0.001$. Note that we choose $m$ and $n$ to be powers of 222 simply for the convenience of implementing randomized Hadamard transform (RHT). In general this is not necessary, since we can still implement it by finding the closest power of 222 larger than $m$ or $n$, enlarging the matrix to that dimension by padding with 0 entries, and truncating back to the original dimension at the end.
@@
[…]
F.1 Testing Projection, Acceleration and Memoization
In this section we test the effect of (i) computing the inner projection steps inexactly, (ii) the adaptive acceleration scheme and (iii) the block memoization technique used in our methods (see Figure 1 and 2 in Section 6, as well as Figure 4 below). For all tasks, we used variants of Kaczmarz++ (Algorithm 5) for solving rectangular linear systems.
To test the the effect of computing the inner projection steps inexactly, we consider two variants of our method:
[…]
  • K++(Cholesky): Kaczmarz++ with exact projections, i.e., Algorithm 5 with function Proj replaced by a direct solve involving $\mathrm{chol}({\mathbf{A}}_{S}{\mathbf{A}}_{S}^{\scriptscriptstyle{\top}}+\lambda{\mathbf{I}})$.
We observed that 888 inner iterations consistently suffices in attaining convergence that nearly matches Kaczmarz++ with exact projections. Thus, for the remaining experiments, we use $t_{\max}=8$ as the default. Next, we consider four variants of our method, as shown in Table 1, to identify the effect of its different components individually. For clarity, we explain their differences below.
@@ F.1 Testing Projection, Acceleration and Memoization
  • Kaczmarz: The classical randomized block Kaczmarz method, without acceleration or block memoization, but still preprocessed with RHT;
@@
  • K++ w/o Accel: Randomized Kaczmarz with block memoization, but without adaptive acceleration, i.e., Algorithm 5 without lines 13-19, and with $\eta=0$ (as opposed to $\eta=\frac{s}{2n}$), meaning that we no longer maintain the adaptive momentum term ${\mathbf{m}}_{t}$ from line 11;
  • K++ w/o Memo: Randomized Kaczmarz with adaptive acceleration, but without block memoization, i.e., Algorithm 5 modified so that in line 6 we always choose to sample a new block $S\sim{[m]\choose s}$, and hence, do not save the Cholesky factors;
@@
  • Full K++: Algorithm 5 as given in the pseudocode, equipped with both adaptive acceleration and block memoization.
Algorithm 5 Kaczmarz++: Kaczmarz type solver for general systems
1:Input: ${\mathbf{A}}\in{\mathds{R}}^{m\times n}$, ${\mathbf{b}}\in{\mathds{R}}^{m}$, block size $s$, iterate ${\mathbf{x}}_{0}$, regularization ${\lambda}$, inner iterations $t_{\max}$, tolerance $\epsilon$;
2:Sample ${\mathbf{D}}\leftarrow\frac{1}{\sqrt{m}}\mathrm{diag}(d_{1},...,d_{m})$ for $d_{i}\sim\mathrm{Rademacher}$;
[…]
4:Initialize ${\mathbf{m}}_{0}\leftarrow\mathbf{0}$, $\rho\leftarrow 0$, $\eta\leftarrow\frac{s}{2n}$, ${\mathcal{B}}\leftarrow\emptyset$, $\zeta\leftarrow\lceil m/s\rceil$, ${\mathcal{E}}_{0},{\mathcal{E}}_{1}\leftarrow 0$, $\tau\leftarrow 2s$;
5:for $t=0,1,...$ do
[…]
7:    ${\mathcal{B}}\leftarrow{\mathcal{B}}\cup\{S\}$ for $S\sim{[m]\choose s}$; $\triangleright$ Sample new subset.
8:  else $S\sim{\mathcal{B}}$; end if
9:  ${\mathbf{r}}_{t}\leftarrow{\mathbf{A}}_{S}{\mathbf{x}}_{t}-{\mathbf{b}}_{S}$$\triangleright$ Use for error estimation.
10:  ${\mathbf{w}}_{t}\leftarrow\mathsf{Proj}({\mathbf{A}},S,{\mathcal{B}},{\mathbf{r}}_{t},\tau,\lambda,t_{\max})$; $\triangleright$ Inner LSQR solver.
11:  ${\mathbf{m}}_{t+1}\leftarrow\frac{1-\rho}{1+\rho}\big{(}{\mathbf{m}}_{t}-{\mathbf{w}}_{t}\big{)}$$\triangleright$ Adaptive momentum.
12:  ${\mathbf{x}}_{t+1}\leftarrow{\mathbf{x}}_{t}-{\mathbf{w}}_{t}+\eta\,{\mathbf{m}}_{t+1}$;
13:  if $t<\zeta\mod 2\zeta$ then ${\mathcal{E}}_{0}\leftarrow{\mathcal{E}}_{0}+\|{\mathbf{r}}_{t}\|^{2}$; else ${\mathcal{E}}_{1}\leftarrow{\mathcal{E}}_{1}+\|{\mathbf{r}}_{t}\|^{2}$;
14:  if $t=2\zeta-1\mod 2\zeta$ then
15:   if ${\mathcal{E}}_{1}\leq\epsilon^{2}\|{\mathbf{b}}\|^{2}$ then return ${\mathbf{x}}_{t+1}$; $\triangleright$ Stopping criterion.
16:   $r\leftarrow ra_{t}+({\mathcal{E}}_{1}/{\mathcal{E}}_{0})(1-a_{t})$; $\triangleright$ Weighted average (5.4).
17:   $\rho\leftarrow 1-r^{1/\zeta}$; $\triangleright$ Convergence rate estimate.
18:   ${\mathcal{E}}_{0},{\mathcal{E}}_{1}\leftarrow 0$;
19:  end if
20:end for
Algorithm 6 Preconditioned LSQR solver $\mathsf{Proj}({\mathbf{A}},S,{\mathcal{B}},{\mathbf{r}},\tau,\lambda,t_{\max})$
1:Input: ${\mathbf{A}}_{S}\in{\mathds{R}}^{s\times n},{\mathbf{r}}\in{\mathds{R}}^{s}$, sketch size $\tau$, regularization $\lambda$, max iterations $t_{\max}$;
2:if $S\not\in{\mathcal{B}}$ then
[…]
5:end if
[…]
8:Return ${\mathbf{w}}\leftarrow{\mathbf{x}}_{1:n}$
[…]
@@
By comparing the curves, we can see that in all cases, adding block memoization slightly worsens the convergence rate (by comparing Kaczmarz and K++ w/o Accel, or K++ w/o Memo and Full K++). This is actually suggested by theory, since with block memoization we are essentially not sampling all possible blocks of coordinates, thus reducing the quality of the block sampling distribution. However, this phenomenon is very insignificant especially when we compare K++ w/o Memo with Full K++, which suggests that our online block selection scheme (Section 5.2) works well, and that $\tilde{O}(\min\{m,n\}/s)$ blocks are sufficient for fast convergence, matching our theory. The computational benefits of block memoization are observed in the FLOPs experiments (Figure 5), once we plot the convergence in terms of arithmetic operations, taking advantage of the saved Cholesky factors.
@@
From the plots, we can also see that the adaptive acceleration plays an important role for both Kaczmarz and Kaczmarz with block memoization, showing a comparable improvement for both cases. We also observe that the effect of adaptive acceleration is more significant when the sketch size $s$ is smaller - this also aligns with the theory, since for smaller $s$ the tail Demmel condition number $\bar{\kappa}_{s}$ is larger, thus the effect of acceleration reducing the dependence on $\bar{\kappa}_{s}$ from second to first power is also more significant.
Figure 4: Convergence plots testing inexact projection steps and different variants of K++ (see Table 1), using two block sizes (continuation of Figure 1 and 2 with effective ranks 25 and 200).
F.2 Comparison with Krylov Subspace Methods
@@
Next, we evaluate the computational cost of our algorithms on kernel matrices based on benchmark machine learning data. Since these are PSD linear systems, we evaluate the convergence of Full CD++, alongside CD++ w/o Memo and CD++ w/o RHT, comparing against Krylov-type methods, conjugate gradient (CG) and GMRES. In this experiment we count the FLOPs it takes to converge for each method, to showcase the advantages of our methods compared to Krylov-type methods.
@@
Figure 5: Computational cost comparison, measuring floating point operations (FLOPs) against the normalized residual error (6.1) for Full CD++, alongside baselines CD++ w/o Memo, CG, and GMRES, and also CD++ w/o RHT (continuation of Figure 3).
@@
For the FLOPs of CG, we look into the source code from scipy.sparse.linalg and approximate it by $2n^{2}+11n$ per iteration. For the FLOPs counting of GMRES, we look into the source code from pyamg.krylov and approximate it by $2n^{2}T+4nT(T+1)$, where $T$ is the number of iterations. For the FLOPs counting of our methods (Full CD++, CD++ w/o Memo, CD++ w/o RHT), we maintain a counter for FLOPs for each iteration. which includes both the cases of using block memoization (Full CD++, CD++ w/o RHT) and not using it (CD++ w/o Memo). Specifically, for the algorithms with block memoization, if a new block is sampled, then we do the Cholesky factorization and increase the FLOPs by $s^{3}/3$, where $s$ is the block size; if we sample from the already sampled set of blocks ${\mathcal{B}}$, then Cholesky factorization is not needed and this cost is omitted. Moreover, for CD++ w/o Memo and Full CD++, there is a pre-processing step of applying the RHT which takes extra FLOPs at the beginning (this is omitted from CD++ w/o RHT). We count them by following Appendix E, and this is reflected in the plots, since the convergence curves of our methods are shifted by the cost of the RHT. Here, we take advantage of our fast SymFHT implementation (Algorithm 2), which reduces that cost by half. Throughout the section, we choose the block size $s=200$, which tends to work well for the size of our test matrices. Further optimizing the block size, or choosing it dynamically, is an interesting direction for future work. We run each algorithm 555 times and take average to reduce the noise.
@@
From the experiments (Figure 5), we can see that CG converges very slowly and is not comparable with other methods. By comparing CD++ w/o Memo and Full CD++, we can see that the block memoization technique gives a significant improvement in FLOPs, since Full CD++ successfully reduces the expensive Cholesky factorization step. This improvement is more significant if we are running more iterations (i.e., aiming for higher accuracy). By comparing Full CD++ and GMRES, we can see that in most cases Full CD++ performs better in the “low to medium accuracy level”, while GMRES sometimes beats Full CD++ in the “high accuracy level”.
@@
For example, for the abalone dataset with Gaussian kernel with width=0.010.010.01, GMRES starts to perform better after the residual reaches $10^{-7}$. However, we note that for abalone, phoneme and california_housing dataset with Gaussian kernel with width=0.10.10.1, or with Laplacian kernel with width=0.010.010.01, our Full CD++ outperforms GMRES even when the residual reaches $10^{-12}$. These phenomena are dependent on the spectrum (especially the top eigenvalues) of matrix ${\mathbf{A}}+\phi{\mathbf{I}}$, which is reflected in the different choices of kernel and width.
Problem CG GMRES Full CD++ CD++ w/o RHT
Dataset Kernel Width 1e-4 1e-8 1e-4 1e-8 1e-4 1e-8 1e-4 1e-8
Abalone Gaussian 0.1 7.63e9 3.06e10 1.47e9 5.17e9 4.64e8 3.26e9 6.68e8 8.97e9
0.01 1.88e9 6.89e9 8.50e8 1.75e9 2.97e8 2.11e9 4.04e8 2.71e9
Laplacian 0.1 8.64e9 3.22e10 2.72e9 8.14e9 2.22e9 8.13e9 1.74e9 6.06e9
0.01 1.11e9 1.37e10 5.41e8 4.11e9 2.40e8 3.09e9 2.20e7 3.06e9
Phoneme Gaussian 0.1 7.80e9 3.29e10 1.79e9 4.98e9 4.86e8 3.23e9 4.61e8 1.00e10
0.01 3.70e8 4.23e9 3.37e8 1.33e9 2.23e8 1.80e9 4.39e6 1.79e9
Laplacian 0.1 7.22e9 3.73e10 2.29e9 8.46e9 1.65e9 8.96e9 1.35e9 1.31e10
0.01 2.49e9 1.41e10 7.81e8 3.97e9 2.80e8 3.10e9 7.47e7 3.01e9
California Housing Gaussian 0.1 2.45e9 3.27e10 1.02e9 6.14e9 2.40e8 3.88e9 3.69e8 9.74e9
0.01 2.15e9 9.51e9 5.75e8 2.18e9 2.27e8 1.99e9 3.64e8 7.13e9
Laplacian 0.1 5.44e9 2.34e10 2.04e9 6.85e9 1.52e9 6.39e9 1.24e9 4.56e9
0.01 1.31e9 1.36e10 5.75e8 4.15e9 2.40e8 3.06e9 4.39e7 3.45e9
Covtype Gaussian 0.1 1.65e10 5.55e10 4.87e9 1.70e10 5.71e9 3.64e10 3.63e9 3.11e10
0.01 9.31e9 3.87e10 2.00e9 7.69e9 9.78e8 5.66e9 1.26e9 1.22e10
Laplacian 0.1 6.79e9 2.24e10 3.23e9 8.34e9 2.85e9 1.03e10 1.88e9 8.61e9
0.01 3.60e9 2.37e10 1.23e9 5.95e9 4.60e8 4.56e9 3.43e8 5.18e9
Synthetic Low-Rank Effective rank =25=25= 25 1.68e9 3.26e9 1.44e9 1.83e9 1.31e9 2.79e9 1.11e9 2.44e9
Effective rank =50=50= 50 2.59e9 5.17e9 2.43e9 3.26e9 1.53e9 3.21e9 1.34e9 2.92e9
Effective rank =100=100= 100 3.16e9 6.89e9 2.65e9 5.67e9 1.91e9 3.89e9 1.94e9 4.16e9
Effective rank =200=200= 200 3.26e9 8.00e9 2.75e9 8.34e9 2.92e9 6.10e9 2.69e9 5.77e9
@@
Table 2: Comparison of the FLOPs needed to achieve given error threshold, $\epsilon\in\{10^{-4},10^{-8}\}$, for different algorithms. Bold values indicate the best performance for a given error threshold. The last column is included to test the effect of RHT preprocessing.
@@
In Table 2 we show the FLOPs it takes for different methods (CG, GMRES, Full CD++, and CD++ w/o RHT) to achieve the given accuracy in detail. Here we set $\epsilon=10^{-4}$ as the mid-level accuracy and $\epsilon=10^{-8}$ as the high-level accuracy. For mid-level accuracy, we can see that Full CD++ outperforms GMRES in 18 out of 20 tasks, showing that our method converges very fast at early stages. For high-level accuracy, we can see that Full CD++ still outperforms GMRES at 14 out of 20 tasks.
Effect of RHT.
We also evaluate the effect of RHT as a preprocessing step (suggested by theory) on the total cost of our algorithm by comparing CD++ w/o RHT and Full CD++. Note that the only difference here is whether or not to use RHT, while the uniform sampling scheme remains the same. By comparing these two curves in Figure 5, we can see that the RHT step itself is very cheap, and can be measured by how much the beginning of the convergence curve of Full CD++ is shifted away from 0 (given that the size of the matrices is $\times 4096$ 4096, the cost of the RHT step is approximately 2.4e8 FLOPs). This cost is almost negligible in most cases, however, it is noteworthy that in some cases (e.g., Abalone dataset with Laplacian kernel with width 0.01), the cost of RHT dominates the cost of the algorithm when converging to mid-level accuracy. Among all the 202020 test matrices (in Figure 3 and 5), RHT provides a significant gain in overall computational performance in 6 cases, and a smaller gain in 3 others. Despite this, even without RHT preprocessing, CD++ w/o RHT still exhibits good convergence for most of the problems and is comparable with GMRES, indicating that our algorithm can be effective even for sparse linear systems, where RHT is not desirable.
@@
To conclude, the experiments in this section show that in the measurement of FLOPs, our CD++ outperforms Krylov methods including CG and GMRES in almost all mid-level accuracy tasks, as well as most high-level accuracy tasks. We also show the feasibility of discarding RHT preprocessing step in our algorithm when the input matrix is sparse.
F.3 Testing Regularization in Projection